# Variational Improvement of the Hartree-Fock Approach to the 2D Hubbard Model

### v1

@@ Abstract
The calculations were carried out on $16\times 16$ lattices with periodic boundary conditions, and our results are shown in the $U-\mu$ plane in Fig. 1, and in the $U-f$ (coupling-density) plane in Fig. 2, where $f=1$ is half-filling. [^1] We have noted that the MATLAB software used here can deliver very slightly different numerical results when the same code is run on different machines with different versions of MATLAB, and this feature has been acknowledged by the company. The points shown in these figures however, d-wave, s-wave, or condensate, are consistent on the two machines we have used. Each of the symbols in Fig. 1 represents a pair of $U,\mu$ values where we have carried out the calculation described above. At the filled square locations, $\mathcal{E}(h)$ for ${q=-1}$, i.e. d-wave, is lower in energy than either $q=1$ (s-wave) or $h=0$ (standard Hartree-Fock). At filled circle locations, $\mathcal{E}(h)$ is lowest for an s-wave condensate. The open square symbols represent locations where standard Hartree-Fock, i.e. $h=0$ with no condensate, have the lowest energy. On average the fractional decrease in energy of the condensate states, as compared to the conventional ($h=0$) states at the same $U,\mu$ values, is 1.3%. The optimum $h$ varies, but is generally $O(10^{-2})$. Qualitatively, from Fig. 2, we see that the condensate disappears completely at low $U<2$ and high $U>10$ values of $U$, and at densities below half-filling. Obviously, with a finite set of $\mu$ values we have not covered the entire $U-\mu$ plane, although nearby filled symbols at a given $U$ probably indicate a continuum of densities with condensate states.
@@
The momentum-space distribution of the condensate is quite different at low and high $U$. An example at $U=2$ and density $0.74$ is shown in Fig. 3. Here the condensate seems to be concentrated mainly at the Fermi surface, moreover the pattern of up/down spins has a stripe pattern, as seen in Fig. 4, which displays the average spin/local magnetization
@@
$=$ $\rho(x,11)+\rho(x,22)-1\ .$
In contrast, the condensate distribution at $U=9$ and density $0.67$ which is displayed in Fig. 5, is much more diffuse, and the pattern of spins (Fig. 6) is a little harder to characterize. But it is clear that in both cases the expectation value of spin densities is not at all space independent. We see that the amplitude of the condensate at small $U=2$, which is concentrated at the Fermi surface, is much larger than the amplitude of the condensate in the more diffuse result at $U=9$. But if one defines a measure of total condensate
$C={1\over L^{2}}\sum_{k}|P(k)|\ ,$ (20)
then the measures at $U=2$ and $U=9$ are comparable, with $C=0.015$ and $C=0.10$ respectively.
Figure 3: Momentum space condensate $P(k)$ at $U=2,\mu=0.1$, where we find density $f=0.74$.
Figure 4: Spin average $D(x)$ on the $16\times 16$ lattice at parameters ${U=2,\mu=0.1}$.
Figure 5: Momentum space condensate $P(k)$ at $U=9,\mu=0.9$, where we find density $f=0.67$.
Figure 6: Spin average $D(x)$ on the $16\times 16$ lattice at parameters ${U=9,\mu=0.9}$.
@@
Since a slight variational improvement of the Hartree-Fock approach yields both s- and d-wave condensates, a natural question is whether we can find a more systematic procedure for obtaining the optimal approximation to the ground state, rather than depending on the choice of a specific condensate-inducing term which, of course, may not be the ideal choice. To improve matters, the general idea is still to combine the Hartree-Fock and variational approaches to optimize a single determinant approximation to the ground state. A maximal (and impractical) procedure would be as follows: On a finite $L\times L$ lattice at fixed $U$ and $\mu$, we apply the Hartree-Fock method to obtain a complete and orthogonal set of of one-particle wave functions $\{\phi_{n}(x,\alpha),n=1,2,...,2L^{2}\}$. The standard Hartree-Fock ground state is shown in (6). Any other Slater determinant can be obtained by replacing the set $\{\phi_{n}\}$ in (6) by a new set $\{\phi^{\prime}_{n}=U_{nm}\phi_{m}\}$, where $U=\exp(i\gamma_{i}T_{i})$ is a $2L^{2}\times 2L^{2}$ unitary matrix which is a member of the $U(N)$ group, with $\{T_{i}\}$ the corresponding generators. Then we may regard the $\gamma_{i}$ as variational parameters, and attempt to vary those parameters in such a way as to minimize the expectation value $\mathcal{E}=\langle\Omega|H|\Omega\rangle$, building $|\Omega\rangle$, as before, out of a subset of the $\{\phi^{\prime}_{n}\}$. In principle, if one could find the set of $\gamma$’s which minimizes $\mathcal{E}$, that would be the best approximation, by a single Slater determinant, to the true ground state of the Hubbard Hamiltonian. In practice there will be a vast number of local minima, but aside from that this “brute force” approach, with the number of variational parameters equal to $4L^{4}$, seems extremely computation intensive. But if it would be the case that the Hartree-Fock approach is only failing close to the Fermi surface, then the problem may be more tractable. For most one-particle states we let $\phi^{\prime}_{n}=\phi_{n}$, and only choose a subset of $K$ one-particle states, in the immediate neighborhood of the Fermi surface, for which we consider $\phi^{\prime}=U\phi$. This time, $U$ is only a $K\times K$ matrix acting on states in the Hilbert space spanned by the $K$ Hartree-Fock one-particle states near the Fermi surface. Of course this still leaves $K^{2}$ variational parameters, but one can imagine some stochastic or relaxation approach which would converge to at least a local minimum. For example, one could cycle through pairs of one particle states in the $K$ subset, and apply at each stage an SU(2) transformation to obtain a new pair of states which are linear combinations of the original pair. These could be accepted, as new one particle states in the Slater determinant, if they lower $\mathcal{E}$. One would of course have to test the sensitivity of the final result to the choice of $K$, to ensure that enough states were chosen.
@@ Acknowledgements.
This research is supported by the U.S. Department of Energy under Grant No. DE-SC0013682.

### v3

@@ IV Variational approach: spatial inhomogeneity
The calculations were carried out on $16\times 16$ lattices with periodic boundary conditions, and our results are shown in the $U-\mu$ plane in Fig. 2, and in the $U-f$ (coupling-density) plane in Fig. 3, where $f=1$ is half-filling. Each of the symbols in Fig. 2 represents a pair of $U,\mu$ values where we have carried out the calculation described above. At the filled square locations, $\mathcal{E}(h)$ for ${q=-1}$, i.e. d-wave, is lower in energy than either $q=1$ (s-wave) or $h=0$ (standard Hartree-Fock). At filled circle locations, $\mathcal{E}(h)$ is lowest for an s-wave condensate. The open square symbols represent locations where standard Hartree-Fock, i.e. $h=0$ with no condensate, have the lowest energy. The optimum $h$ varies, but is generally $O(10^{-2})$. Qualitatively, from Fig. 3, we see that the condensate disappears completely at low $U<4$, and at densities below half-filling. Obviously, with a finite set of $\mu$ values we have not covered the entire $U-\mu$ plane, although nearby filled symbols at a given $U$ probably indicate a continuum of densities with condensate states. We have carried out the calculation up to $U=12$.
@@
Figure 4: Optimal values of the variational parameter $h$ in the $U-\mu$ plane for the state of lowest energy, at points where we find either s- or d-wave condensates.
The optimal values of $h>0$ in the $U-\mu$ plane, for the $U,\mu$ points which have been studied (see Fig. 2) are shown in Fig. 4. Zero values, in this and in Figs. 5 and 6, are either points in which the unmodified Hartree-Fock state has the lowest energy, or else were not investigated. Coordinates on the $x,y$ axes of these bar plots are required, by the plotting software, to be integer valued, and for this reason we have multiplied our $\mu$ values in this and the following two plots by a factor of ten. Figure 5 displays the energy difference between the energy $E_{0}$ of the $h=0$ unmodified state, and the energy $E_{c}$ of the d- or s-wave state, whichever is lower in energy, when that energy is below $E_{0}$. The difference between the d-wave $E_{d}$ and s-wave $E_{s}$ energies, when both are less than $E_{0}$ are shown in Fig. 6. These energy differences are usually smaller, sometimes much smaller, than $E_{0}-E_{c}$.
Figure 5: The difference $E_{0}-E_{c}$ between the $h=0$ energy $E_{0}$, and the energy $E_{c}$ of the optimal condensate ground state, at points where we find $E_{c}<E_{0}$.
One awkward feature of Fig. 3 is the fact that points of d- and s-wave condensate are interspersed, and often next to one another in the $U-$density plane. On the other hand, the energies of the optimal d-wave and s-wave states are often so close together, differing by a fraction of one percent, that a change in lattice volume can reverse the energy ordering of the two states. We have seen this in a number of points we have checked, whose d- and s-wave energies differ by O($10^{-4}$). In such cases we find that a point of s-wave condensate on a $16\times 16$ lattice may convert, at the same density, to a d-wave condensate on a $24\times 24$ lattice, and vice versa. Therefore only two statements can be made with any confidence: condensates tend to lower the energy (compared to the Hartree-Fock ground state) at higher densities, generally at $\mu>1,f>0.6$ with $U>4$, and spatial inhomogeneity is a necessary but not sufficient condition for their appearance.
Figure 6: The magnitude of the difference between the optimal s- and d-wave ground state energies, when those energies are both less than the corresponding $E_{0}$.
Figure 7: Momentum space condensate $P(k)$ at $U=8,\mu=1.4,h=0.01$, where we find density $f=0.83$.
The momentum-space distribution of the condensate is quite different from the concentration at the Fermi surface seen in Fig. 1. An example at $U=8,\mu=1.4,h=0.01$ and density $0.83$ which is displayed in Fig. 7, is much more diffuse. In this case the $\mathcal{E}(h)=-1.802$, while $\mathcal{E}(0)=-1.789$. The maximum amplitude of the condensate at small $U=2$ in Fig. 1, which is concentrated at the Fermi surface, is much larger (by about a factor of forty) than the maximum amplitude of the condensate in the more diffuse result at $U=8$ shown in Fig. 7. But if one defines a measure of average condensate magnitude
$C={1\over L^{2}}\sum_{k}|P(k)|\ ,$ (57)
then the measures at $U=2$ and $U=8$ are $C=0.010$ and $C=0.0033$ respectively, differing only by a factor of three. This is simply because, although the maximum amplitudes at $U=8$ are small, the condensate is more diffuse in $k$-space as compared with the concentration at the Fermi surface at $U=2$, and there are small but significant contributions to $C$ over a greater momentum range.
The average spin at each site is given by
@@
$=$ $\rho(x,11)+\rho(x,22)-1\ ,$
and the result for parameters $U=8,\mu=1.4,h=0.01$ is shown in Fig. 8. It is clear in this case, as in all cases where we have found $\mathcal{E}(h)<\mathcal{E}(0)$ for some $h>0$, that the expectation value of spin densities is far from uniform. Defining an average spin magnitude per site,
$m_{av}={1\over L^{2}}\sum_{x}|D(x)|\ ,$ (59)
we find that in all cases where $\mathcal{E}(h)<\mathcal{E}(0)$ for some $h>0$, $m_{av}$ ranges from 0.30 to 0.94. Averaging over all such cases, the value is 0.63 with a standard deviation of 0.15 in this set. Since $N_{\uparrow}=N_{\downarrow}$, the spatial average of $D(x)$ must vanish. Then the fact the spatial average $D(x)$ is zero yet the magnitude at each site deviates quite strongly from zero implies spatial inhomogeneity of some kind, as in Fig. 8. In contrast, running the numerical Hartree Fock program at the parameters $U=2,\mu=0.1$ as in Fig. 1, where the $h=0$ state is lowest energy, we find $m_{av}=1.6\times 10^{-16}$, i.e. near perfect spatial homogeneity.
Figure 8: Spin average $D(x)$ on the $16\times 16$ lattice at parameters $U=8,\mu=1.4,{h=0.01}$. $x$ and $y$ axes correspond to $x,y$ coordinates on the lattice, and the color of each square indicates the value of $D(x)$ at that position.
V Discusson and generalization
We have found that a slight variational improvement of the Hartree-Fock approach lowers the ground state energy, as compared to the unimproved state, at sufficiently large density and coupling, and that these lower energy states are associated with either s- or d-wave condensates. Only spatially inhomogeneous states, with spin expectation values $D(x)$ significantly different from zero, seem to have this property. For a review of the possible relation between stripes and superconductivity, see Tranquada_2020 .
@@ V Discusson and generalization
A natural question is whether we can find a more systematic procedure for obtaining the optimal approximation to the ground state, rather than depending on the choice of a specific condensate-inducing term which, of course, may not be the ideal choice. To improve matters, the general idea is still to combine the Hartree-Fock and variational approaches to optimize a single determinant approximation to the ground state. A maximal (and impractical) procedure would be as follows: On a finite $L\times L$ lattice at fixed $U$ and $\mu$, we apply the Hartree-Fock method to obtain a complete and orthogonal set of of one-particle wave functions $\{\phi_{n}(x,\alpha),n=1,2,...,2L^{2}\}$. The standard Hartree-Fock ground state is shown in (7). Any other Slater determinant can be obtained by replacing the set $\{\phi_{n}\}$ in (7) by a new set $\{\phi^{\prime}_{n}={\cal U}_{nm}\phi_{m}\}$, where ${\cal U}=\exp[i\gamma_{i}T_{i}]$ is a $2L^{2}\times 2L^{2}$ unitary matrix which is a member of the $U(N)$ group, with $\{T_{i}\}$ the corresponding generators. Then we may regard the $\gamma_{i}$ as variational parameters, and attempt to vary those parameters in such a way as to minimize the expectation value $\mathcal{E}=\langle\Omega|H|\Omega\rangle$, building $|\Omega\rangle$, as before, out of a subset of the $\{\phi^{\prime}_{n}\}$. In principle, if one could find the set of $\gamma$’s which minimizes $\mathcal{E}$, that would be the best approximation, by a single Slater determinant, to the true ground state of the Hubbard Hamiltonian. In practice there will be a vast number of local minima, but aside from that this “brute force” approach, with the number of variational parameters equal to $2L^{2}$, seems extremely computation intensive. But if it would be the case that the Hartree-Fock approach is only failing close to the Fermi surface, then the problem may be more tractable. For most one-particle states we let $\phi^{\prime}_{n}=\phi_{n}$, and only choose a subset of $K$ one-particle states, in the immediate neighborhood of the Fermi surface, for which we consider $\phi^{\prime}={\cal U}\phi$. This time, ${\cal U}$ is only a $K\times K$ matrix acting on states in the Hilbert space spanned by the $K$ Hartree-Fock one-particle states near the Fermi surface. Of course this still leaves $K^{2}$ variational parameters, but one can imagine some stochastic or relaxation approach which would converge to at least a local minimum. For example, one could cycle through pairs of one particle states in the $K$ subset, and apply at each stage an SU(2) transformation to obtain a new pair of states which are linear combinations of the original pair. These could be accepted, as new one particle states in the Slater determinant, if they lower $\mathcal{E}$. One would of course have to test the sensitivity of the final result to the choice of $K$, to ensure that enough states were chosen.
@@ Acknowledgements.
We thank Richard Scalettar for helpful correspondence. This research is supported by the U.S. Department of Energy under Grant No. DE-SC0013682.